OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (3, 4) and has a slope of 2? y = x +

OpenStudy (anonymous):

Ok so this is simple. the equation is y = mx + b m being slope and b being the y-intercept do you understand so far??

OpenStudy (anonymous):

( the y-intercept is when -- on a graph -- the line that connects the points going through the y-axis. Heres an example..)

OpenStudy (anonymous):

 this would be the y-intercept

OpenStudy (anonymous):

Since we already know the slope (2), we can place that into the sport where m was so it was y = mx + b but now its y = 2x + b

OpenStudy (anonymous):

The next step is replacing b with an actual number.so using the coordinates (3,4), we can easily find the y-intercept (b).

OpenStudy (anonymous):

How do we find it?

OpenStudy (anonymous):

XD haha! Ok so do you know what the y-intercept is??

OpenStudy (anonymous):

or do you know what the meaning is??

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i know what y- intercept is :)

OpenStudy (anonymous):

ok so you know that the full point slope equation is y = mx +b right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

So youre just having trouble identifying where to place the numbers?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Ok so in the equation y = mx + b, m is the slope. So using the information you gave me, i know that the slope is 2. so now the equation is...

OpenStudy (anonymous):

(Youre turn XD)

OpenStudy (anonymous):

let me seee

OpenStudy (anonymous):

wait what is the answer tho?

OpenStudy (anonymous):

Well OpenStudy doesn't really like when people give out answers. Walking you through it will only take a second with cooperation :)

OpenStudy (anonymous):

All you have to do is find the y intercept which is very easy

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

So we know the slope, yeah? Its two. The variable m, is representing slope... so what is the equation now...?